# BroadGen: A Framework for Generating Effective and Efficient Advertiser Broad Match Keyphrase Recommendations

### v2

@@ 4.4.1. Dependency on Input Queries
Table 4. Effect of number of pre-queries on BroadGen.
4.4.2. Number of Keyphrases
Achieving a good balance with recall and precision for high F1 is difficult due to the limitation on the number of broad match recommendations (see Section 1.1.2) — fewer keyphrases lead to fewer queries being reached. BroadGen(Aug) also maintains balance when increasing the number of broad match recommendations. This is evident from Figure 5 where the recall as well as precision increase with the number of recommendations. The trend is notably elevated for CAT_4, as its ambiguous characteristics lead to item diversity, resulting in a sparse distribution of relevant queries.
4.5. Online Deployment
4.5.1. Production Implementation
The daily recommendations of BroadGen(Aug) is made possible through eBay’s PySpark-based (Foundation, 2025) batch inference system. Unlike GraphEx (Mishra et al., 2025) where a daily differential batch inference can be implemented, it is not required for our model due to its nature of taking Query as an input, which is constantly changing over time. However, the speed of our model is so fast, that with 1000 Spark executors and 4 cores, it only takes 45 minutes to finish the whole batch inference on a space of 1.5 billion items and their corresponding Queries. The inputs of the production model are pre-queries with at least 5 impressions in the past 30 days augmented with exact keyphrases that pass the Bert relevance filter (Dey et al., 2025a) from two of our recall models — fastText (Bojanowski et al., 2017; Joulin et al., 2017, 2016) and Graphite (Mishra et al., 2024). These two recall models are trained as Extreme Multi-Label Classifiers (Agrawal et al., 2013) with relevant past queries as labels for exact keyphrase recommendations. With BroadGen’s processing ability, we can incorporate newer queries on a daily basis accounting for shifting label distributions (see Section 1.1.1) thus solving Challenge 1.2.4.
4.5.2. Impact
@@
5. Conclusion and Future Work
We introduce a new framework for advancing match type recommendations, addressing the challenges associated with exact keyphrase suggestions in sponsored search advertising. Our distributed model, BroadGen, employs a string clustering approach to offer broad match recommendations, efficiently handling up to a billion items in under an hour. Without access to ground truth match type keyphrases, we construct keyphrases by modeling the underlying token correspondence in historical queries. We also propose a unique evaluation methodology and metrics to evaluate broad match performance when real-world impressions and clicks are not available. In the future, we aim to expand our approach by implementing neural network equivalents to the cores of our framework. Such as, using text embeddings for input matrix generation while using a k-means or similar clustering algorithm for grouping the queries. Text summarization for the cluster representation core can be tricky as they tend to generate too generic or specific keyphrases. With the adoption of more complex cores based on LLMs we can subsequently cater to broader definitions of semantic broad match (eBay Inc., 2025; Inc., 2025; Broder et al., 2007) and extend beyond token-based broad match in the future. Training different aspects of such a model will be difficult due to the absence of concrete supervised signals and the constraints of broad/phrase match recommendation. Moreover, the latency and infrastructure costs of these models will also be a concern due to growing number of items and queries.

### v3

@@ IV-D1 Dependency on Input Queries
TABLE IV: Effect of number of pre-queries on BroadGen.
F1 (RRE) PTR
Model - Clustering algorithm CAT_1 CAT_2 CAT_3 CAT_4 CAT_1 CAT_2 CAT_3 CAT_4
BroadGen(Aug) - Affinity Propagation 0.39 0.49 0.42 0.69 0.20 0.24 0.22 0.28
BroadGen(Aug) - Bayesian Gaussian Mixture 0.36 0.45 0.43 0.64 0.18 0.20 0.18 0.25
BroadGen(Aug) - Agglomerative 0.42 0.53 0.47 0.67 0.25 0.32 0.28 0.36
TABLE V: Comparing F1 (RRE) and PTR scores of BroadGen(Aug), when the clustering core is replaced with different algorithms. $\alpha=1.0$ and $\beta=1.5$ was used for PTR.
Figure 5: Comparison of Precision, Recall and F1 for BroadGen(Aug) across keyword counts and categories.
IV-D2 Various Clustering Algorithms
Different clustering algorithms can replace the BroadGen clustering core outlined in Section III-C. Algorithms like HDBSCAN [46] are unsuitable for BroadGen due to its requirement for a minimum cluster size, which can lead to grouping dissimilar queries. Table V compares BroadGen’s performance when its agglomerative core is replaced with Affinity Propagation [47] and Bayesian Gaussian Mixture [48]. We present F1 scores based on the RRE metric alongside PTR scores. The Agglomerative method shows superior F1 scores in CAT_1-3, albeit slightly lower in CAT_4. Its PTR scores are notably higher than the other algorithms, as it offers greater flexibility in clustering, enabling extraction of longer keyphrases beneficial for match-type keyword recommendations. KMeans and Spectral Clustering were also evaluated, but their excessive runtime of over 140140ms and 190190ms per input on average impedes daily batch processing.
IV-D3 Number of Keyphrases
Achieving a good balance with recall and precision for high F1 is difficult due to the limitation on the number of broad match recommendations (see Section I-A2) — fewer keyphrases lead to fewer queries being reached. BroadGen(Aug) also maintains balance when increasing the number of broad match recommendations. This is evident from Figure 5 where the recall as well as precision increase with the number of recommendations. The trend is notably elevated for CAT_4, as its ambiguous characteristics lead to item diversity, resulting in a sparse distribution of relevant queries.
IV-E Online Deployment
IV-E1 Production Implementation
The daily recommendations of BroadGen(Aug) is made possible through eBay’s PySpark-based [49] batch inference system. The speed of our model is considerable — with 1000 Spark executors and 4 cores, it only takes 45 minutes to finish the whole batch inference on a space of 1.5 billion items and their corresponding Queries. The inputs of the production model are pre-queries with at least 5 impressions in the past 30 days augmented with exact keyphrases that pass the Bert relevance filter [12] from two of our recall models — fastText [50, 51] and Graphite [3]. These two recall models are trained as Extreme Multi-Label Classifiers [52] with relevant past queries as labels for exact keyphrase recommendations. With BroadGen’s processing ability, we can incorporate newer queries on a daily basis accounting for shifting label distributions (see Section I-A1) thus solving Challenge I-B4.
IV-E2 Impact
@@
IV-E3 Sensitivity Analysis
As noted in Section I-A1, exact-match keyphrases are vulnerable to changes in buyer search behavior, while broad-match keyphrases offer some adaptability by utilizing a flexible matching mechanism. BroadGen(Aug) advances this strategy by incorporating past queries to adjust to shifting query distributions. To evaluate its performance, we monitored BroadGen over 30 consecutive days post-advertising campaign initiation. Figure 6 illustrates that BroadGen keyphrases consistently improve beyond the initial 10 days, encountering a minor dip after three weeks, but ultimately achieving an average daily performance increase of 0.46%. This indicates enhanced effectiveness and stability against query drift. Conversely, in the control set, exact-match and broad-match keyphrases both exhibit a continual decline, though broad-match keyphrases fare slightly better, declining at –1.02% daily, compared to –1.11% for exact-match keyphrases.
Figure 6: Performance drift over time for the BroadGen keyphrases in the treatment group vs baseline exact match and broad match keyphrases from the control group
V Conclusion and Future Work
We introduce a new framework for advancing non-exact match type recommendations, addressing the challenges associated with exact keyphrase suggestions in sponsored search advertising. Our distributed model, BroadGen, employs a string clustering approach to offer broad match recommendations, efficiently handling up to a billion items in under an hour. Without access to ground truth match type keyphrases, we construct keyphrases by modeling the underlying token correspondence in historical queries. We also propose a unique evaluation methodology and metrics to evaluate broad match performance when real-world impressions and clicks are not available. In the future, we aim to expand our approach by implementing neural network equivalents to the cores of our framework. Such as, using text embeddings for input matrix generation while using a k-means or similar clustering algorithm for grouping the queries. Text summarization for the cluster representation core can be tricky as they tend to generate too generic or specific keyphrases. With the adoption of more complex cores based on LLMs we can subsequently cater to broader definitions of semantic broad match [26, 27, 53] and extend beyond token-based broad match in the future. Training different aspects of such a model will be difficult due to the absence of concrete supervised signals and the constraints of broad/phrase match recommendation. Moreover, the latency and infrastructure costs of these models will also be a concern due to growing number of items and queries.